Hockey teams receive 2 points for a win and 1 point for a tie. The wild cats once one a championship with 60 points. They won 9 more games than they tied. How many wins and how many ties did the Wildcats have? I thought I set this problem up correctly but it didnt work out right??: Hockey teams receive 2 points for a win and 1 point for a tie. The wild cats once one a championship with 60 points. They won 9 more games than they tied. How many wins and how many ties did the Wildcats have? I thought I set this problem up correctly but it didnt work out right??
Answer by josmiceli(2035)

Let w = number of games won
Let t = number of games tied
w = t + 9
As for the points:
2*w + 1*t = 60
Use the expression for w in the 1st equation for w in
the 2nd equation
2*(t + 9) + t = 60
2t + 18 + t = 60
3t = 42
t = 14
and since
w = t + 9
w = 14 + 9
w = 23
There were 23 wins and 14 ties
check:
2*w + 1*t = 60
2*23 + 1*14 = 60
46 + 14 = 60
60 = 60
OK